The speed of light C is related to the permeability and thr permittivity e by the expression
C^2*uo*e = 1
(i) Show that this equation is homogenous
(ii) Calculate the value of uo
1
Expert's answer
2021-01-29T20:02:10-0500
  1. "[c^2] = \frac{m^2}{s^2}", "[\mu_0] = \frac{H}{m}, [\epsilon_0] = \frac{F}{m}", thus "[\mu_0 \epsilon_0] = \frac{F \cdot H}{m^2} = \frac{\frac{F}{C^2} \cdot HA^2 s^2}{m^2}", as "C=A\cdot s", but "C^2\/F = HA^2 = J" and thus "[\mu_0 \epsilon_0]= \frac{s^2}{m^2}, [c^2\mu_0 \epsilon_0]=1" and so this equation is homogeneous.
  2. From this identity we find "\mu_0 = \frac{1}{\epsilon_0 c^2} \approx \frac{1}{8.85\cdot 10^{-12}\cdot 9\cdot 10^{16}} = 1.26 \cdot 10^{-6} H/m". Commonly the value of "\mu_0" is approached by "4\pi \cdot 10^{-7} H/m".